Given f(x) = 2x+3 and g(x) = 4-5x, determine the x-value for which f(2x) = g(1-x).
Answer by amarjeeth123(569)

We are given that f(x)=2x+3 and g(x)=4-5x
We have to determine the x-value for which f(2x)=g(1-x)
f(2x)=2(2x)+3=4x+3
g(1-x)=4-5(1-x)=4-5+5x=5x-1
Plugging in the values we get,
4x+3=5x-1
5x-4x=3+1
x=4.
The value of x is 4.